# On the Sylvester program and Cayley algorithm for vector partition reduction

### v3

@@ 5 Vector partition with collinear columns
where $u_{j}$ are positive integers and vector ${\bf C}$ represents the GCD vector that divides elementwise all the collinear vectors ${\bf c}_{i},1\leq i\leq n<m$. The problem (62) is equivalent to
@@
where $K$ is computed using only positive $C_{r}>0$. Introduce a vector ${\bf u}=\{u_{1},u_{2},\ldots,u_{n}\}$ and a competent matrix with $(m-n)$ columns $\bar{\bf D}=\{{\bf c}_{n+1},\ldots,{\bf c}_{m}\},$ for which the corresponding partition $W({\bf S}(k),\bar{\bf D})$ admits the reduction either as the sum of $W_{i}^{l}$ or the mixture of $W^{l}_{i}$ and $\bar{W}^{l}_{i}$. The number of solutions of the system (63) for given value of $k$ is equal to a number of solutions of the first equation given by the partition $W({\bf S}(k),\bar{\bf D})$ multiplied by a number of solutions of the second equation, i.e., by a scalar partition $W(k,{\bf u})$. Then the vector partition $W({\bf s},{\bf D})$ is equivalent to a convolution
@@
This approach can be extended to the case when the matrix ${\bf D}$ has $r>1$ groups of the collinear columns. Each $p$-th group is characterized by its GCD vector ${\bf C}_{p}$ and the $n_{p}$ integers $u_{p,i}$ making up the vector ${\bf u}_{p}=\{u_{p,1},u_{p,2},\ldots,u_{p,n_{p}}\}$. Introduce the vector ${\bf k}=\{k_{1},k_{2},\ldots,k_{r}\}$ and (63) turns into
@@
Consider a special case $n=m-1$, i.e., when the competent matrix $\bar{\bf D}=\{{\bf c}\}$ has a single column ${\bf c}$. The vector partition $W({\bf S}({\bf k}),\{{\bf c}\})=1$ only when ${\bf S}({\bf k})=k_{0}{\bf c}$ for some nonnegative integer $k_{0}$, otherwise it vanishes. Thus we observe that $W({\bf S}({\bf k}),\bar{\bf D})=\delta_{{\bf S}({\bf k}),k_{0}{\bf c}}$ and we obtain using (66)
@@
The extreme case arises for $n=m$, when each column of the matrix ${\bf D}$ belongs to one of $r$ groups of the collinear columns. Use ${\bf k}^{\prime}=\{k_{1},k_{2},\ldots,k_{r-1}\}$ to write
@@
The example of partitions with collinear columns is discussed in Appendix C where we describe the reduction procedure for the matrix used in computation of the Kronecker coefficients [9, 10].
@@ 6 Iterative reduction of vector to scalar partitions
The basic step of iterative reduction procedure of the vector partition corresponding to the $l$-row matrix ${\bf E}$ described in Section 4 includes elimination of the last row and a single $i$-th column of ${\bf E}$. It results in a single partition based on the $(l-1)$-row matrix for the P-column or a superposition of partitions built on similar matrices for the NP-column. Each such matrix undergoes similar transformation in the consecutive iterative steps leading to a combination of scalar partitions. The total number of iterations required to reach this stage is $l-1$.
@@
How can one define the partition function in this case? The last row ${\bf R}_{n}$ of the matrix corresponds to a scalar partition defined by (1) that requires to find all sets ${\bf x}_{K}$ of nonnegative integer solutions. Assume that we find all such sets, their number is given by a scalar partition $W(S_{n},\{R_{n1},\ldots,R_{np}\})$. Now substitute each solution ${\bf x}_{K}$ into the remaining $n-1$ equations and retain only those that satisfy them all. The number of such solutions is the vector partition $W({\bf E}_{\nu})$. This definition fails when the last row has zero entries as the number of solutions ${\bf x}_{K}$ in this case is infinite. Nevertheless, one still can proceed with the evaluation procedure in these cases, this problem is addressed in Section 4.3 where we prove that the contribution of a column with the last zero element is zero.
@@ 7 Conclusive remarks
In this manuscript we present an implementation of the Sylvester program aimed to reduction of vector (compound) partition to scalar partitions. The first step in this direction was made in 1860 by Cayley [4] when he showed that the double partition problem subject to a set of conditions admits a compact solution in which the $i$-th column ${\bf c}_{i}$ of a positive generator matrix ${\bf D}$ leads to a single scalar partition contribution $W_{i}^{2}$ (see Section 3.1). The author of this manuscript have suggested an alternative approach that solves the problem when the original Cayley method fails because the conditions are not satisfied [14] .
The results presented in the manuscript show that Sylvester strongly underestimated the number of scalar partitions required for computation of the vector partition – their minimal number in case of $l\times m$ positive matrix might significantly surpass $m!/(m-l+1)!$ which in its turn is $(l-1)!$ times larger than predicted by Sylvester in [17]. We address the case of collinear columns of the generator matrix in Section 5 and show that in this case a vector partition boils down to a convolution of scalar partitions.
@@
All known approaches to evaluation of vector partitions (see [1, 15, 3, 18]), their later modifications [2, 5, 8] and the software packages based of these methods successfully determine the chambers but fail to produce the explicit expressions for the vector partitions as they replace the original generating function by the signed sum of the simpler partial generating functions valid in specific chamber. This means that for each new argument column ${\bf s}$ one has to perform the Taylor expansion of the partial generating functions to extract the desired coefficient. Our method is the first one that permits the derivation of a single explicit formula for the vector partition which is valid for an arbitrary value of ${\bf s}$. The usage of this formula does not require the explicit knowledge of the chambers but nevertheless it implicitly contains their description.
@@ Appendix A Triple partition reduction example
${\bf E}=\left(\begin{array}[]{ccccc}s_{1}&1&2&1&1\\ s_{2}&2&1&3&4\\ s_{3}&3&2&3&2\end{array}\right),$ (A1)
@@
${\bf E}_{1}=\left(\begin{array}[]{cccc}3s_{1}-s_{3}&4&0&1\\ 3s_{2}-2s_{3}&-1&3&8\end{array}\right),\quad{\bf E}_{2}=\left(\begin{array}[]{cccc}3s_{1}-s_{3}&0&4&1\\ 3s_{2}-3s_{3}&-3&-3&6\end{array}\right),$ (A7)
${\bf E}_{3}=\left(\begin{array}[]{cccc}2s_{1}-s_{3}&-1&2&-1\\ 2s_{2}-4s_{3}&-8&-6&-6\end{array}\right),\quad{\bf E}_{4}=\left(\begin{array}[]{cccc}2s_{1}-2s_{3}&-4&-4&-2\\ 2s_{2}-s_{3}&1&3&6\end{array}\right),$ (A12)
@@
$\tilde{\bf E}_{1}=\left(\begin{array}[]{cccc}3s_{1}-s_{3}+4&-4&0&1\\ 3s_{2}-2s_{3}-1&1&3&8\end{array}\right),$ $\tilde{\bf E}_{2}=\left(\begin{array}[]{cccc}3s_{1}-s_{3}+4&0&-4&1\\ s_{2}-s_{3}-2&1&1&2\end{array}\right),$ (A18)
$\tilde{\bf E}_{3}=\left(\begin{array}[]{cccc}2s_{1}-s_{3}&1&-2&1\\ s_{2}-2s_{3}-10&4&3&3\end{array}\right),$ $\tilde{\bf E}_{4}=\left(\begin{array}[]{cccc}s_{1}-s_{3}&-2&-2&-1\\ 2s_{2}-s_{3}&1&3&6\end{array}\right).$ (A23)
@@ Appendix B Alternative triple partition reduction
${\bf E}^{\prime}=\left(\begin{array}[]{ccccc}s_{2}&2&1&3&4\\ s_{1}&1&2&1&1\\ s_{3}&3&2&3&2\end{array}\right),$ (B1)
@@
$\tilde{\bf E}^{\prime}_{1}=\left(\begin{array}[]{cccc}2s_{2}-s_{3}&-1&-3&-6\\ s_{1}-s_{3}-5&2&2&1\end{array}\right),\quad\tilde{\bf E}^{\prime}_{2}=\left(\begin{array}[]{cccc}3s_{2}-2s_{3}&-1&8&3\\ 3s_{1}-s_{3}&4&1&0\end{array}\right),$ (B7)
$\tilde{\bf E}^{\prime}_{3}=\left(\begin{array}[]{cccc}s_{2}-s_{3}-1&1&-1&2\\ 3s_{1}-s_{3}&0&4&1\end{array}\right),\quad\tilde{\bf E}^{\prime}_{4}=\left(\begin{array}[]{cccc}s_{2}-2s_{3}-7&-3&4&3\\ 2s_{1}-s_{3}-2&2&1&1\end{array}\right).$ (B12)
@@
The second reduction step turns each double partition $W(\tilde{\bf E}^{\prime}_{k})$ into a sum of scalar partitions, but this time the contribution of the columns $\{3,0\}^{T}$ in $\tilde{\bf E}^{\prime}_{2}$ and $\{1,0\}^{T}$ in $\tilde{\bf E}^{\prime}_{3}$ vanishes. Thus we find
@@
Appendix C Partition with collinear columns

### v4

@@ 5.1 Partition convolution superposition
where $u_{j}$ are positive integers and one can always select the vector $\bf C$} to be a P-column.
The problem (62) is equivalent to
@@
where $K$ is computed using only positive $C_{r}>0$. Introduce a vector ${\bf u}=\{u_{1},u_{2},\ldots,u_{n}\}$ and a competent matrix with (mn)(m-n) columns $\bar{\bf D}=\{{\bf c}_{n+1},\ldots,{\bf c}_{m}\},$ for which the corresponding partition $W({\bf S}(k),\bar{\bf D})$ admits the reduction either as the sum of $W_{i}^{l}$ or the mixture of $W^{l}_{i}$ and $\bar{W}^{l}_{i}$. The number of solutions of the system (63) for given value of $k$ is equal to a number of solutions of the first equation given by the partition $W({\bf S}(k),\bar{\bf D})$ multiplied by a number of solutions of the second equation, i.e., by a scalar partition $W(k,{\bf u})$. Then the vector partition $W({\bf s},{\bf D})$ is equivalent to a convolution
@@
This approach can be extended to the case when the matrix $\bf D$} has $r>1$ groups of the C-columns. Each $p$-th group is characterized by the P-column ${\bf C}_{p}$ and the $n_{p}$ integers $u_{p,i}$ making up the vector ${\bf u}_{p}=\{u_{p,1},u_{p,2},\ldots,u_{p,n_{p}}\}$. Introduce the vector ${\bf k}=\{k_{1},k_{2},\ldots,k_{r}\}$ and (63) turns into
@@
Consider a special case n=m1n=m-1, i.e., when the competent matrix $\bar{\bf D}=\{{\bf c}\}$ has a single column $\bf c$}. The vector partition $W({\bf S}({\bf k}),\{{\bf c}\})=1$ only when ${\bf S}({\bf k})=k_{0}{\bf c}$ for some nonnegative integer $k_{0}$, otherwise it vanishes. Thus we observe that $W({\bf S}({\bf k}),\bar{\bf D})=\delta_{{\bf S}({\bf k}),k_{0}{\bf c}}$ and we obtain using (66)
@@
The extreme case arises for $n=m$, when each column of the matrix $\bf D$} belongs to one of $r$ groups of the C-columns. Use ${\bf k}^{\prime}=\{k_{1},k_{2},\ldots,k_{r-1}\}$ to write
@@
The example of partitions with C-columns is discussed in Appendix C where we describe the reduction procedure for the matrix used in computation of the Kronecker coefficients [10, 11].
5.2 Collinear columns elimination
An alternative approach to the reduction of the incompetent matrix was introduced by Sylvester in the fifth lecture of his course on the partitions [20] where he suggested to supply the system ${\bf s}={\bf D}\cdot{\bf x}$ with a new unknown $x_{0}$ and a single auxiliary equation to generate a new system $\tilde{\bf s}=\tilde{\bf D}\cdot\tilde{\bf x}$ with $\tilde{\bf x}=\{x_{0},x_{1},x_{2},\ldots,x_{m}\}$. The $(l+1)\times(m+1)$ integer competent matrix $\tilde{\bf D}=\{\tilde{\bf c}_{0},\tilde{\bf c}_{1},\tilde{\bf c}_{2},\ldots,\tilde{\bf c}_{m}\}$ consists of $(m+1)$ columns $\tilde{\bf c}_{i}=\{c_{i0},c_{i1},c_{i2},\ldots,c_{il}\}^{T}$. The elements $c_{i0}$ and the value of $s_{0}$ in the added equation must be chosen to satisfy two conditions – (a) each nonnegative integer solution of the original system ${\bf s}={\bf D}\cdot{\bf x}$ should correspond to a single solution of the system $\tilde{\bf s}=\tilde{\bf D}\cdot\tilde{\bf x}$, and (b) the matrix $\tilde{\bf D}$ should be void of C- and NP-columns. The first condition is equivalent to
$W(\tilde{\bf s},\tilde{\bf D})=W({\bf s},{\bf D}),$ (70)
[…]
The choice of the added equation is not unique and following the Sylvester advice we use
$s_{0}=x_{0}-\sum_{i=1}^{m}\delta_{i}x_{i},$
[…]
$\tilde{\bf D}=\left[\begin{array}[]{ccccc}1&-\delta_{1}&-\delta_{2}&\ldots&-\delta_{m}\\ 0&c_{11}&c_{21}&\ldots&c_{m1}\\ 0&c_{12}&c_{22}&\ldots&c_{m2}\\ \ldots&\ldots&\ldots&\ldots\\ 0&c_{1l}&c_{2l}&\ldots&c_{ml}\\ \end{array}\right],\quad\quad\tilde{\bf s}=\left[\begin{array}[]{c}s_{0}\\ s_{1}\\ s_{2}\\ \ldots\\ s_{m}\end{array}\right].$ (71)
Each integer nonnegative solution ${\bf x}=\bm{\xi}=\{\xi_{1},\xi_{2},\ldots,\xi_{m}\}$ of the system ${\bf s}={\bf D}\cdot{\bf x}$ corresponds to the following solution $\tilde{\bf x}$ of the extended system
$x_{i}=\xi_{i},\ (1\leq i\leq m),\quad\quad x_{0}=\sum_{i=1}^{m}\delta_{i}\xi_{i},$ (72)
where we set $s_{0}=0$. We observe that the condition (70) is satisfied as for each solution $\bm{\xi}$ of the original Diophantine equation $s={\bf d}\cdot{\bf x}$ the system $\tilde{\bf s}=\tilde{\bf D}\cdot\tilde{\bf x}$ has exactly one solution (72).
The selection of the positive $\delta_{i}$ values is a (partially heuristic) process governed by the requirement that all columns of the matrix $\tilde{\bf D}$ should have $\mbox{gcd}(\tilde{\bf c}_{i})=1$ and be not collinear to each other. To satisfy the first condition one has to choose $\delta_{i}$ that does not divide the GCD of the column ${\bf c}_{i}$ of the original matrix $\bf D$}, i.e., $\delta_{i}\nmid\mbox{gcd}({\bf c}_{i})$. As for the second condition one has to consider several cases. Start from the matrix $\bf D$} having a single group of $n$ C-columns ${\bf c}_{j}=u_{j}{\bf C}$ with integer $u_{j}$ and $\bf C$} chosen to be a P-column. To make the $\tilde{\bf c}_{j}$ noncollinear P-columns use the following simple rules
$\mbox{gcd}(u_{j},\delta_{j})=1,\quad\delta_{j}/u_{j}\neq\delta_{k}/u_{k},\quad 1\leq j\neq k\leq n.$ (73)
When the matrix $\bf D$} has several groups of the C-columns, the columns belonging to the different groups are not collinear to each other and thus the procedure of $\delta_{i}$ selection can be done using the rules (73) independently within each group. The search of $\delta_{i}$ is discussed in the illustrative example presented in Appendix D.
It should be underlined that the rules (73) leave a lot of freedom in the $\delta_{i}$ selection so that the matrix $\tilde{\bf D}$ used for the further reduction is not unique. The reduction itself is done by using the Cayley elimination algorithm that produces $(m+1)$ vector partitions $W(\tilde{\bf s}^{i},\tilde{\bf D}^{i})$, but $W(\tilde{\bf s}^{0},\tilde{\bf D}^{0})$ evaluates to zero as explained in Section 4.3 and we end up with just $m$ vector partitions.
From comparison of the matrices $\bf D$} and $\tilde{\bf D}$ it follows that the elements of the vector ${\bf s}^{1}$ given by (45) and the components of $\tilde{\bf s}^{1}$ from the second to the last one coincide – $s^{1}_{i}=\tilde{s}^{1}_{i+1},\ 1\leq i\leq l$. The same is true for the columns ${\bf c}^{1}_{i}$ and $\tilde{\bf c}^{1}_{i}$ with $2\leq i\leq m$. We already noted in Section 4 for the matrix $\bf D$} with two C-columns ${\bf c}_{1}=u_{1}{\bf C}$ and ${\bf c}_{k}=u_{k}{\bf C}$ after reduction we obtain for the column ${\bf c}^{1}_{k}={\bf 0}$. This means that in the column $\tilde{\bf c}^{1}_{k}$ the first element $\tilde{c}^{1}_{k1}=(u_{k}\delta_{1}-u_{1}\delta_{k})C_{m}$ is nonzero while all other elements vanish, and the contribution of $\tilde{\bf c}^{1}_{k}$ in the next reduction round equals zero.
5.3 Comparison of the methods
In the convolution method in Section 5.1 the transfer of a group of the C-columns from the matrix $\bf D$} to the argument $\bf S$} diminishes the number of columns in $\bar{\bf D}$ so that it may significantly decrease the number of the produced scalar partitions. On the other hand, this relatively small number of the scalar partitions enter the summands in the multiple sums that for the large values of $\bf s$} may lead to very large number of terms requiring evaluation.
Turning to the alternative matrix extension method in Section 5.2 we note that it replaces a single $l\times m$ matrix by $m$ matrices of the same size that might significantly increase the total number of the scalar partitions in the final expression. The same time this larger number of scalar partitions are evaluated only once to obtain the value of the original vector partition. For a given vector partition this algorithm also allows to generate multiple equivalent superpositions of scalar partitions.
@@ 6 Iterative reduction of vector to scalar partitions
The basic step of iterative reduction procedure of the vector partition corresponding to the $l$-row matrix $\bf E$} described in Section 4 includes elimination of the last row and a single $i$-th column of $\bf E$}. It results in a single partition based on the (l1)(l-1)-row matrix for the P-column or a superposition of partitions built on similar matrices for the NP-column. Each such matrix undergoes similar transformation in the consecutive iterative steps leading to a combination of scalar partitions. The total number of iterations required to reach this stage is l1l-1.
@@
How can one define the partition function in this case? The last row ${\bf R}_{n}$ of the matrix corresponds to a scalar partition defined by (1) that requires to find all sets ${\bf x}=\bm{\xi}_{K}$ of nonnegative integer solutions. Assume that we find all such sets, their number is given by a scalar partition $W(S_{n},\{R_{n1},\ldots,R_{np}\})$. Now substitute each solution $\bm{\xi}_{K}$ into the remaining n1n-1 equations and retain only those that satisfy them all. The number of such solutions is the vector partition $W({\bf E}_{\nu})$. This definition fails when the last row has zero entries as the number of solutions $\bm{\xi}_{K}$ in this case is infinite. Nevertheless, one still can proceed with the evaluation procedure in these cases, this problem is addressed in Section 4.3 where we prove that the contribution of a column with the last zero element is zero.
@@ 7 Conclusive remarks
In this manuscript we present an implementation of the Sylvester program aimed to reduction of vector (compound) partition to scalar partitions [19, 20]. The first step in this direction was made in 1860 by Cayley [5] when he showed that the double partition problem subject to a set of conditions admits a compact solution in which the $i$-th column ${\bf c}_{i}$ of a positive generator matrix $\bf D$} leads to a single scalar partition contribution $W_{i}^{2}$ (see Section 3.1). The author of this manuscript have suggested an alternative approach that solves the problem when the original Cayley method fails because the conditions are not satisfied [15] .
The results presented in the manuscript indicate that Sylvester strongly underestimated the number of scalar partitions required for computation of the vector partition – their minimal number in case of $l\times m$ positive matrix might significantly surpass m!/(ml+1)!m!/(m-l+1)! which in its turn is (l1)!(l-1)! times larger than predicted by Sylvester in [19]. We address the case of collinear columns in the matrix $\bf D$} in Section 5 and show that a vector partition either boils down to a convolution of scalar partitions or it is replaced by an equivalent one with the larger matrix having no such columns.
@@
All known approaches to evaluation of vector partitions (see [1, 17, 3, 21]), their later modifications [2, 6, 9] and the software packages based of these methods successfully determine the chambers but fail to produce the explicit expressions for the vector partitions as they replace the original generating function by the signed sum of the simpler partial generating functions valid in specific chambers. This means that for each new argument column $\bf s$} one has to perform the Taylor expansion of the partial generating functions to extract the desired coefficient. Our method is the first one that permits the derivation of a single explicit formula for the vector partition which is valid for an arbitrary value of $\bf s$}. The usage of this formula does not require the explicit knowledge of the chambers but nevertheless it implicitly contains their description.
@@ Appendix A Triple partition reduction example
${\bf E}=\left[\begin{array}[]{ccccc}s_{1}&1&2&1&1\\ s_{2}&2&1&3&4\\ s_{3}&3&2&3&2\end{array}\right],$ (A1)
@@
${\bf E}_{1}=\left[\begin{array}[]{cccc}3s_{1}-s_{3}&4&0&1\\ 3s_{2}-2s_{3}&-1&3&8\end{array}\right],\quad{\bf E}_{2}=\left[\begin{array}[]{cccc}3s_{1}-s_{3}&0&4&1\\ 3s_{2}-3s_{3}&-3&-3&6\end{array}\right],$ (A7)
${\bf E}_{3}=\left[\begin{array}[]{cccc}2s_{1}-s_{3}&-1&2&-1\\ 2s_{2}-4s_{3}&-8&-6&-6\end{array}\right],\quad{\bf E}_{4}=\left[\begin{array}[]{cccc}2s_{1}-2s_{3}&-4&-4&-2\\ 2s_{2}-s_{3}&1&3&6\end{array}\right],$ (A12)
@@
$\tilde{\bf E}_{1}=\left[\begin{array}[]{cccc}3s_{1}-s_{3}+4&-4&0&1\\ 3s_{2}-2s_{3}-1&1&3&8\end{array}\right],$ $\tilde{\bf E}_{2}=\left[\begin{array}[]{cccc}3s_{1}-s_{3}+4&0&-4&1\\ s_{2}-s_{3}-2&1&1&2\end{array}\right],$ (A18)
$\tilde{\bf E}_{3}=\left[\begin{array}[]{cccc}2s_{1}-s_{3}&1&-2&1\\ s_{2}-2s_{3}-10&4&3&3\end{array}\right],$ $\tilde{\bf E}_{4}=\left[\begin{array}[]{cccc}s_{1}-s_{3}&-2&-2&-1\\ 2s_{2}-s_{3}&1&3&6\end{array}\right].$ (A23)
@@ Appendix B Alternative triple partition reduction
${\bf E}^{\prime}=\left[\begin{array}[]{ccccc}s_{2}&2&1&3&4\\ s_{1}&1&2&1&1\\ s_{3}&3&2&3&2\end{array}\right],$ (B1)
@@
$\tilde{\bf E}^{\prime}_{1}=\left[\begin{array}[]{cccc}2s_{2}-s_{3}&-1&-3&-6\\ s_{1}-s_{3}-5&2&2&1\end{array}\right],\quad\tilde{\bf E}^{\prime}_{2}=\left[\begin{array}[]{cccc}3s_{2}-2s_{3}&-1&8&3\\ 3s_{1}-s_{3}&4&1&0\end{array}\right],$ (B7)
$\tilde{\bf E}^{\prime}_{3}=\left[\begin{array}[]{cccc}s_{2}-s_{3}-1&1&-1&2\\ 3s_{1}-s_{3}&0&4&1\end{array}\right],\quad\tilde{\bf E}^{\prime}_{4}=\left[\begin{array}[]{cccc}s_{2}-2s_{3}-7&-3&4&3\\ 2s_{1}-s_{3}-2&2&1&1\end{array}\right].$ (B12)
@@
The second reduction step turns each double partition $W(\tilde{\bf E}^{\prime}_{k})$ into a sum of scalar partitions, but this time the contribution of the LZ-columns $\{3,0\}^{T}$ in $\tilde{\bf E}^{\prime}_{2}$ and $\{1,0\}^{T}$ in $\tilde{\bf E}^{\prime}_{3}$ vanishes and we find
@@
Appendix C Partition with collinear columns – convolution algorithm